# Deep Learning of Multivariate Extremes via a Geometric Representation

### v1

@@ B.9 Algorithm for computing the extended ADF diagnostic
Consider a sample $\{\bm{x}_{j}\}_{j=1}^{n}$ of independent copies of $\bm{X}$, with $\{(r_{j},\bm{w}_{j})\}_{j=1}^{n}$ denoting the corresponding radial-angular observations, and let $q$ denote quantile level close to 1. The extended ADF diagnostic is computed using Algorithm 1.
@@
         $t_{\bm{w}_{j}}^{k}\leftarrow\min_{i=1,\ldots,d}\{x_{k,i}/w_{j,i}\}$
     end for
     $\mathbf{t}_{\bm{w}_{j}}\leftarrow(t_{\bm{w}_{j}}^{1},\ldots,t_{\bm{w}_{j}}^{n})$
     $\hat{u}_{\bm{w}_{j}}\leftarrow\texttt{quantile}(\mathbf{t}_{\bm{w}_{j}},q)$
     if $r_{j}>\hat{u}_{\bm{w}_{j}}$ then
         $e_{b}\leftarrow\hat{\Lambda}(\bm{w}_{j})(r_{j}-\hat{u}_{\bm{w}_{j}})$
@@
Note that $r_{j}=t_{\bm{w}_{j}}^{j}$; hence, the radial component can be considered an observation of the min-projection for the corresponding angle. We compare the resulting quantiles, $(e_{1},\ldots,e_{b})^{T}$, to the corresponding theoretical quantiles using a QQ plot.
@@ Appendix C Supplementary figures and tables
Figure A1: Plot illustrating the result described in Proposition 2.8. Here, the red line denotes the unit-level set $\partial\mathcal{G}$ of a Gaussian copula with $\rho=-0.5$. Moreover, the blue and cyan points, and regions, illustrate the re-scaling to obtain $\tilde{r}_{\bm{w}}$ for the angle $\bm{w}=(0.8,-\sqrt{1-0.8^{2}})$. Setting $\bm{c}=\text{sgn}(\bm{w})=(1,-1)$, the green dotted line denotes the unit-level set $\partial\mathcal{G}$ of $\bm{c}\bm{X}$, with the corresponding scaling procedure for the angle $\bm{c}\bm{w}=(0.8,\sqrt{1-0.8^{2}})$ illustrated by the orange points and regions. One can observe that the rescaling procedures are analogous in both orthants.
@@
Figure A3: Goodness-of-fit diagnostics for location 01: Truncated gamma QQ plot (left), the extended ADF diagnostic (centre), and return level probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A4: Goodness-of-fit diagnostics for location 46: Truncated gamma QQ plot (left), the extended ADF diagnostic (centre), and return level probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A5: Goodness-of-fit diagnostics for location 85: Truncated gamma QQ plot (left), the extended ADF diagnostic (centre), and return level probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A6: The estimated unit-level set (left) and the extended ADF (right) for hs, ws, and mslp at location 46 of the metocean data. Data for the three variables are plotted on standard Laplace margins.
Figure A7: The estimated unit-level set (left) and the extended ADF (right) for hs, ws, and mslp at location 85 of the metocean data. Data for the three variables are plotted on standard Laplace margins.
Figure A8: Pairwise distributions of ws, hs, and mslp from the metocean data at location 46, transformed into common Laplace margins; the red shapes outline the corresponding estimated bivariate unit-level set slices.
Figure A9: Pairwise distributions of ws, hs, and mslp from the metocean data at location 85, transformed into common Laplace margins; the red shapes outline the corresponding estimated bivariate unit-level set slices.
Figure A10: Goodness-of-fit diagnostics for hs: Truncated gamma QQ plot (left), the extended ADF diagnostic (centre), and return level probability estimates (right) for the DeepGauge model fitted to wave height data at 5 locations.

### v2

@@ B.9 Algorithm for computing the extended ADF diagnostic
Consider a sample $\{\bm{x}_{j}\}_{j=1}^{n}$ of independent copies of $\bm{X}$, and let $q$ denote quantile level close to 1. Given any angle $\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$, the extended ADF diagnostic is computed using Algorithm 1. We then compare the resulting quantiles, $(e_{1},\ldots,e_{b})^{T}$, to the corresponding theoretical quantiles using a QQ plot.
@@
for $k\leftarrow 1$ to $n$ do
     $t_{\bm{w}}^{k}\leftarrow\min_{i=1,\ldots,d}\{x_{k,i}/w_{i}\}$
end for
$\mathbf{t}_{\bm{w}}\leftarrow(t_{\bm{w}}^{1},\ldots,t_{\bm{w}}^{n})$
$\hat{u}_{\bm{w}}\leftarrow\texttt{quantile}(\mathbf{t}_{\bm{w}},q)$
@@
     if $t_{\bm{w}}^{k}>\hat{u}_{\bm{w}}$ then
         $e_{b}\leftarrow\hat{\Lambda}(\bm{w})(t_{\bm{w}}^{k}-\hat{u}_{\bm{w}})$
@@ Appendix C Supplementary figures and tables
Figure A1: Plot illustrating the result described in Proposition 2.8. Here, the red line denotes the unit-level set $\partial\mathcal{G}$ of a Gaussian copula with $\rho=-0.5$. Moreover, the blue and cyan points, and regions, illustrate the re-scaling to obtain $\tilde{\mathfrak{r}}_{\bm{w}}$ for the angle $\bm{w}=(0.8,-\sqrt{1-0.8^{2}})$. Setting $\bm{c}=\text{sgn}(\bm{w})=(1,-1)$, the green dashed line denotes the unit-level set $\partial\mathcal{G}$ of $\bm{c}\bm{X}$, with the corresponding scaling procedure for the angle $\bm{c}\bm{w}=(0.8,\sqrt{1-0.8^{2}})$ illustrated by the orange points and regions. One can observe that the rescaling procedures are analogous in both orthants.
@@
Figure A3: Goodness-of-fit diagnostics for location 01: Truncated gamma QQ plot (left), the extended ADF diagnostic at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$ (centre), and return level set probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A4: Goodness-of-fit diagnostics for location 46: Truncated gamma QQ plot (left), the extended ADF diagnostic at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$ (centre), and return level set probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A5: Goodness-of-fit diagnostics for location 85: Truncated gamma QQ plot (left), the extended ADF diagnostic at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$ (centre), and return level set probability estimates (right) for the DeepGauge model fitted to hs, ws, and mslp.
Figure A6: The estimated unit-level (left) and extended ADF (right) sets for hs, ws, and mslp at location 46 (on standard Laplace margins)
Figure A7: The estimated unit-level (left) and extended ADF (right) sets for hs, ws, and mslp at location 85 (on standard Laplace margins)
Figure A8: Scaled bivariate sample clouds with $\epsilon=0.01$ for pairwise combinations of ws, hs, and mslp at location 46 on standard Laplace margins; the red lines describe the corresponding estimated bivariate unit-level set slices.
Figure A9: Scaled bivariate sample clouds with $\epsilon=0.01$ for pairwise combinations of ws, hs, and mslp at location 85 on standard Laplace margins; the red lines describe the corresponding estimated bivariate unit-level set slices.
Figure A10: Bootstrapped log-probability estimates for each hypercube at location $01$. The colouring of each boxplot corresponds to the regions in Figure 10.
Figure A11: Bootstrapped log-probability estimates for each hypercube at location $46$. The colouring of each boxplot corresponds to the regions in Figure 10.
Figure A12: Bootstrapped log-probability estimates for each hypercube at location $85$. The colouring of each boxplot corresponds to the regions in Figure 10.
Figure A13: Goodness-of-fit diagnostics for hs: Truncated gamma QQ plot (left), the extended ADF diagnostic at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$ (centre), and return level set probability estimates (right) for the DeepGauge model fitted to wave height data at five locations.
Appendix D Case study with eight dimensional transect
To further demonstrate the efficacy of the DeepGauge framework, we also apply the model to wave height data across $d=8$ locations in the transect; these locations are illustrates in Figure A14. We remark that very few approaches for multivariate extremes are available to dimensions of this size (Lederer and Oesting,, 2023).
Figure A14: Locations used to study severe ocean events. Locations where we study the joint extremal dependence of ws, hs, and mslp are circled in blue. Locations for studying the joint extremal dependence in hs are highlighted in black.
The diagnostics for the $d=8$ case are illustrated in Figures A15 and A16. One can observe generally good agreement across all metrics, indicating the model fits are reasonable. We observe that in a similar manner to the $d=5$ case, there is slightly more divergence at extremely high levels of the quantile plots compared to the $d=3$ examples.
Figure A15: Scaled bivariate sample clouds with $\epsilon=0.015$ for three pairwise combinations of the hs variable over the eight locations on standard Laplace margins; the red lines describe the estimated bivariate unit-level set slices.
Figure A16: Goodness-of-fit diagnostics for hs: Truncated gamma QQ plot (left), the extended ADF diagnostic at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$ (centre), and return level set probability estimates (right) for the DeepGauge model fitted to wave height data at eight locations.